### v1

# The Translation Trap Hypothesis for Multilingual Generation with Large Language Models The Translation Trap Hypothesis: Translation as a Constraint on Multilingual Generation

Abstract
Multilingual generation with large language models (LLMs) is often of poor quality for mid- to low-resource languages. Building on insights from interpretability, we demonstrate the existence of an implicit task-solvingtranslation pipeline for generation, whereby the model first solves the required task in a largely target-language-agnostic manner, and subsequently translates answer concepts into the intended target language. We hypothesize that the failure of the translation stage is an important culprit for the observed low quality of final outputs, and formalize this as the translation barrier hypothesis. We test this hypothesis for a word translation task across language pairs, using logit lens to observe model processing in intermediate layers. We find that a significant portion of overall failures indeed stems from translation failure, or the model’s inability to translate correctly solved intermediate concepts into the target language. This is especially true for low-resource target languages. Our results highlight an important hurdle for end-to-end multilingual generation, and lend guiding insights for future work seeking to improve multilinguality in LLMs.111github.com/JHU-CLSP/translation-barrier
The Translation Barrier Hypothesis:
Multilingual Generation with Large Language Models Suffers from Implicit Translation Failure
Niyati Bafna1, Tianjian Li1, Kenton Murray1, David R. Mortensen2, David Yarowsky1, Hale Sirin1, and Daniel Khashabi1 1Johns Hopkins University, Center for Language and Speech Processing; 2Language Technologies Institute, Carnegie Mellon University {nbafna1,tli104}@jhu.edu
1 Introduction
Large language model (LLM) generation in mid-to-low-resource languages (LRLs) is of notoriously worse quality than that in high-resource languages (HRLs) (Robinson et al., 2023; Hendy et al., 2023; Cahyawijaya et al., 2024; Jiao et al., 2023; Shen et al., 2024), and further suffers the problem of off-targetness, where an LM fails to output text in the intended language of generation (Zhang et al., 2020; Li and Murray, 2023; Marchisio et al., 2024). While ongoing work attempts to use explicit cascaded approaches, i.e. LLM generation into HRLs, followed by translation into the intended target, for multilingual generation (Shi et al., 2022; Liu et al., 2024), or attempt to include support for more languages in LLMs for end-to-end generation, we still lack a systematic understanding of the reasons for and mechanisms of the above failures.
Work from the field of interpretability shows that intermediate layer representations of multilingual LLMs are close to English representations even when prompted to generate in other languages, and that the models demonstrate language-specific behavior only in late layers of the model (Wendler et al., 2024; Foroutan et al., 2022; Tang et al., 2024; Kojima et al., 2024). Building upon this intuition, we posit a model-internal task-solvingtranslation cascade, with the middle layers of the LLM responsible for task-solving, or discovering the required output concepts, in a target-language-agnostic manner, and the last few layers responsible for realizing those concepts in the target language, through translation.
We present the translation barrier hypothesis, illustrated and stated in Figure 1, regarding the poor quality of multilingual generation. This hypothesis has important consequences for end-to-end multilingual generation. If translation failure, as opposed to task-solving failure, is the dominant culprit for the generation gap, then we can modularize or investigate architectural alternatives for the final LLM layers, or seek to induce better transfer in other ways, while still relying on powerful LLM task processing for end-to-end generation in LRLs. On the other hand, if task-solving capabilities themselves fail for LRL targets, explicit cascading methods become more viable options, calling for more investment in external specialized HRL-LRL MT systems.
Our work explores the translation barrier on a word translation task, for two mainstream multilingual models. We choose word translation as a minimal generation task: in this simple case, task-solving consists simply of realizing the semantics or the content of the input term, whereas translation refers to generating the target language term for that concept. We use logit lens to retrieve model-internal hypotheses across different layers, assessing whether the model has arrived at the appropriate semantics for a particular input regardless of language; we also assess final on-target accuracy at the final layer. This allows us to demonstrate the existence of the posited pipeline. We then test our hypothesis, by quantifying the extent to which either stage in the pipeline is responsible for final failure. We make the following contributions:
We introduce the translation barrier hypothesis, highlighting a key mechanism of failure for multilingual generation with LLMs.
Working with a word translation task, we provide evidence for the existence of an implicit task-solvingtranslation pipeline for LLM generation over (source) (target) language pairs, developing a framework within which we can assess and compare the contribution of task-solving and translation failure to total final failure.
We characterize the languages in which task-solving occurs, going beyond English-centric previous work to show that intermediate layer representations are in fact multilingual to some extent and dominated by model-supported languages, with English foremost.
We find that translation failure dominates () final failure for and of all language pairs for aya-23 and llama-3.1 respectively for our task, and that it especially forms the bottleneck for generation in low-resource or unsupported target languages. On the other hand, we find that task-solving is the bottleneck in the case of a low-resource source language.
2 Experimental Setup
Word translation dataset
Given a word translation task between a source and target language, we are interested in judging the semantics of intermediate layer outputs regardless of language. We create a multiparallel dataset consisting of translations of words aligned across all languages of interest. We do this by taking source words in English, split equally among nouns, verbs, adjectives, and adverbs, and using Google Translate API222translate.google.com to collect a set of translations per word, including synonyms, for all languages and for English itself. This allows us to construct a lexicon between any source and target language for the same concepts, as well as assess off-target accuracy in all other languages per concept. We manually inspected the source words and lexicons for correctness and coverage (see Appendix B for more details).
Models
We choose two performant mainstream models with a differing extents of multilinguality to conduct our experiments on: Aya-23-8B (aya-23) (Aryabumi et al., 2024) and Llama-3.1-8B-Instruct (llama-3.1) (Grattafiori et al., 2024). These are both 8-billion-parameter, decoder-only Transformer models, fine-tuned for instruction-following in and languages respectively. See Appendix C for generation hyperparameters and our prompting setup.
Languages
By virtue of their data-hungry nature, mainstream LLMs only support a limited set of usually high- and mid-resource languages, but may be used and are claimed to demonstrate incident multilingualism in other languages (Blevins and Zettlemoyer, 2022). We study target languages, including all languages supported by either model,333We count simplified and traditional Chinese as separate; thus we say that aya-23 supports languages. Thai is the only language supported by llama-3.1 but not aya-23. and a mix of additional mid- and low-resource languages unsupported by both models. See Appendix A for the full list of supported and unsupported languages. We use three different source languages against all the above targets: a high-resource supported language (Spanish), a mid-resource supported language (Hindi), and a mid-to-low resource unsupported language (Telugu). Note that “source” language refers to that of the translation task input; the task instruction is always provided in English. We consider all the above languages as candidates for task-solving languages in intermediate model layers.
3 Quantifying Task-solving and Translation failure
The translation barrier hypothesis presupposes the existence of a model-internal task-solvingtranslation pipeline (validated and discussed in § 4), with model success predicated on task-solving success followed by translation success. We can study this hypothesis in the context of a given task by quantifying the extent to which translation failure is responsible for poor final outputs, as opposed to task-solving failure.
Obtaining intermediate layer outputs
We use an interpretability tool called logit lens (Nostalgebraist, 2020) to obtain text outputs from intermediate layers. Let be an input prompt and be the output of layer . Logit lens applies the unembedding matrix and softmax to to produce a probability distribution over the vocabulary at layer :
where norm refers to RMSNorm Zhang and Sennrich (2019) for llama-3.1 and Layer Normalization (Ba et al., 2016) for aya-23. We apply greedy decoding with this distribution to produce text output from the intermediate layer .
Logit lens only operates at the token level; however, terms in our dataset may be multi-token. We iteratively append the most probable token of the final layer output (standard greedy decoding) to and apply a logit lens at every decoding step, enabling us to decode multi-token outputs from intermediate layers.
Defining translation loss
Given a metric which measures the quality of final model outputs for our task, suppose that we can measure the quality of task-solving at intermediate layers with a comparable metric . We then define translation loss for a given source word and reference y,
where is the text output at model layer , and is the total number of layers. Aggregating instance level over the dataset yields the dataset level translation loss:
The first term in Equation 1 measures task-solving success: we consider the task-solving performance at the best intermediate (non-final) layer for a given input. The second term measures final success at output layer , i.e. the success of the task-solvingtranslation pipeline. Thus , the difference between these, measures performance loss due to failed (internal) translation.
Defining for word translation
Our word translation dataset contains a set of equivalents per concept (source word ) per target language . The metric for measuring final accuracy is a binary exact match metric with any of the target language reference translations. We consider the model to have task-solved successfully at a given layer if the layer output expresses the semantics of the input in any non-source language. Thus,
where is the set of all languages and is the source language. Given that the binary accuracy metric straightforwardly allows us to compute total failure as the number of source words with failed final translations , we can now compute translation loss proportion ():
Note that in this setup, the complement gives us task-solving failure as a fraction of total failure. Thus, computing allows us to compare the proportion of blame that either stage of the pipeline bears for total final failure.
Layerwise analysis
For each layer output, we are interested in its task-solving accuracy, measured by the exact match metric , and its language. We use our dataset language tag for the layer output in case of a match against our listed language equivalents. For inaccurate outputs, we perform language ID on the layer output with NLLB language ID model (NLLB Team et al., 2022).444huggingface.co/facebook/fasttext-language-identification/ We only retain the tag if it belongs to our set of target languages, which includes all languages supported by the model i.e. plausible intermediate languages, and discard it if not. This is because intermediate layer output text may be gibberish or malformed, in which case we expect the language ID model to return a noisy random tag. See Appendix E for a manual evaluation of the reliability of the resulting tags. We track accuracy and language of output for the last ten layers of each model.555We inspected all layers with preliminary experiments and found that intermediate accuracy emerges in a model- and language- dependent manner around this layer. Including earlier intermediate layers yielded similar results for our experiments.
4 Results and Discussion
4.1 Visualizing the pipeline
In Figure 2, we plot on-target and off-target accuracy for four language pairs over all outputs for which we found a reliable language tag. As targets, we include French and Indonesian as representative HRL targets, and Marathi as an example of an unsupported LRL target. We also include all three source languages. We find the patterns discussed below hold generally for all language pairs; see Appendix F for more examples.
We first observe the effect of the source language. While both models show high off-target accuracy with Hindi and Spanish as sources, aya-23 behaves differently with respect to Telugu: regardless of target, it yields low intermediate or final accuracy. This indicates aya-23 simply fails to comprehend Telugu task inputs, resulting in failed task-solving even for HRL targets such as French. This is less true of llama-3.1, which shows similar patterns with Telugu as source as with Hindi and Spanish.
We then look at the effect of the target language. We consistently observe high off-targetness in middle layers, including for Spanish and other supported or high-resource languages, with considerable off-target accuracy, which subsequently decreases as the model pivots to the desired language. For supported or HRL targets, there is a corresponding increase in on-target outputs and accuracy in later layers, indicating successful translation. Note that while HRLs like Indonesian may not be supported by the LLM (llama-3.1), we still see similar patterns for these languages as for supported HRLs, indicating that language resourcedness and presumably its presence in pre-training corpora play an important role in this regard. On the other hand, for Marathi and other unsupported or low-resource languages, we simply observe a total drop in correct (on-target or off-target) intermediate outputs as the model moves away from off-target equivalents but also fails to translate correctly, resulting in poor final outputs. This provides intuition for the translation barrier for LRL generation with LLMs. We note that languages pairs with English as a target behave differently from the above, with on-targetness and accuracy emerging early and continuing until the final layer. This is not surprising, given the dominance of English in model training data.
The consistent nature of these plots, with off-target accuracy in middle layers (successfully or unsuccessfully) converted to on-target accuracy, can be interpreted as the model first “task-solving”, or realizing correct answer concepts in some language(s), and then converting or translating those to the intended language. We find that and of all on-target final correct answers for all language pairs are also identified as correct at some intermediate layer in a non-target language for llama-3.1 and aya-23 respectively (see Appendix F for language-wise breakdowns). This, along with the above visualizations, shows evidence of a task-solvingtranslation pipeline in the model. This also validates our technique of observing intermediate accuracy.
4.2 Characterizing task-solving languages
The above observations raise the question of what language(s) task-solving occurs in, referred to task-solving languages, and whether these differ by language pair. In order to investigate this, we first identify the task-solving layers of the model by locating the switch between the task-solving and translation stages of the generation pipeline, and then study the distribution of languages in the task-solving layers.
Layer of switch
We use the proportion of the target language among all accurate layer outputs to track the switch between the two stages, since this will be high in the translation stage, and plot this by layer for five representative languages in Figure 3 for both models.666We also look at the proportion of target language outputs regardless of accuracy by layer; this results in similar trends. See Appendix H for complete plots.
In general, both models show a significance increase in target language presence around the last four layers. We imagine that this is model- and task- dependent. We also compute the specific layer with the maximum increase in target language presence from the previous layer ranges from layer, and find that this ranges to over different language pairs (see Appendix G for details). We found no correlation between the layer of switch and accuracy at the final layer over all language pairs.
Distribution over task-solving languages
As per the above, we consider all layers up to the fourth-last layer as task-solving layers,777This is a conservative but heuristic approximation: as shown in Figure 3, the switch between the stages is fairly fluid around the last few layers. and look at the distribution over languages for all identified intermediate correct (off-target) answers in these layers. We show the aggregate distribution over all language pairs in Figure 4 for both models.
We see that English dominates the intermediate representations for both models, although considerably more so for llama-3.1. This is expected, given that aya-23 is much more multilingual than llama-3.1. However, while previous work such as (Wendler et al., 2024; Etxaniz et al., 2024; Schut et al., 2025) focus exclusively on English as the intermediate representation language, we find that it only accounts for and of correct intermediate layer outputs for aya-23 and llama-3.1 respectively, with the rest of the distribution spread over other high-resource languages, with a tail of LRLs.888Note that we only compute this distribution over our considered languages. These include all model-supported languages: we observe that these constitute the bulk of the probability mass. Our experiments indicate that including other unsupported languages will largely only modify the long tail of the distribution. While this shows that multilingual LLMs are undeniably English-centric, it also demonstrates some extent of multilinguality, and raises questions about what information might be represented in other languages.
We therefore also look at this distribution conditioned on target language (averaged over source languages), and find that it looks similar at middle layers, showing an increase of target language presence near final layers. (See Appendix H for details.) Thus, we find that the task-solving stage of our demonstrated pipeline occurs in an English-dominant mix of supported and high-resource languages, and that this mix of languages is largely agnostic of the target language in middle layers.
4.3 Quantifying the translation barrier
Figure 2 suggests that the model is often able to solve the task, but fails to generate / translate correctly into LRLs. We quantify this effect by computing: (a) task-solving success or intermediate accuracy (the first term in Equation 1): the percentage of task outputs that were correct at some intermediate layer in some language (b) final success (the second term in Equation 1): on-target accuracy at the final layer, and (c) , as per Equation 2: the proportion of total failure that can be attributed to translation failure.
First, we look at these three quantities averaged over all target languages for each source language in Table 1. We see that intermediate accuracy is much higher than final accuracy in all cases, with much less variance, indicating that the former is stabler and less dependent on the target language than the latter. This ties in with our understanding of the relatively target-language agnostic task-solving step, followed by a target-language dependent translation step. Overall, translation loss accounts for a high percentage of total failure (more than in all cases except one). this lends evidence to the translation barrier hypothesis.
| Src. | Final Acc. | Int. Acc. | TLP | |
| aya | spa* | 42.0 28.1 | 84.6 7.2 | 71.6 7.3 |
| -23 | hin* | 37.4 24.3 | 76.1 7.2 | 59.6 8.4 |
| tel | 9.5 13.0 | 25.3 11.3 | 17.5 4.0 | |
| Avg | 29.6 | 62.0 | 49.6 | |
| llama | spa* | 39.8 29.4 | 84.2 9.8 | 68.0 18.3 |
| -3.1 | hin* | 33.7 23.8 | 75.2 7.8 | 58.2 15.3 |
| tel | 31.6 21.1 | 69.4 7.8 | 52.1 12.6 | |
| Avg. | 35.0 | 76.3 | 59.4 |
We also see that final and intermediate accuracy are roughly similar for both models, with the exception of for Telugu, for which aya-23 is much worse. This is perhaps surprising given than aya-23 is more multilingual than llama-3.1, supporting as opposed to just of our target languages.
Next, we look at intermediate accuracy, final accuracy, and by language pair and model to understand how these vary in more detail. We show these results for target languages against each source language for both models in Figure 5, with the complete results for all target languages in Appendix J. Broadly, high intermediate accuracy indicates successful task-solving; if it is accompanied by high final accuracy, that indicates a subsequently successful translation stage. As discussed in § 3, is a function of these quantities, measuring the proportion of failure cases (i.e. the complement of final accuracy) that can be attributed to translation loss (representing cases that were correctly task-solved but incorrectly translated).
Role of translation failure
We find that translation failure dominates total failure () for and of all language pairs for and aya-23 and llama-3.1 respectively. This provides direct evidence for the translation barrier hypothesis for individual language pairs.
We further find distinct patterns for high-resource versus low-resource target languages. We observe that intermediate accuracy stays high even for low-resource unsupported target languages (like Nepali and Hebrew) for Hindi and Spanish as sources: in line with our previous observations, this indicates that the model is able to correctly task-solve even for low-resource targets. In contrast, and as expected, we observe that final accuracy is high for supported target HRLs (like German, Italian, and Portuguese, for both models) but drops drastically for low-resource target languages (like Tamil and Swahili), some of which are supported, such as for Greek with aya-23. High intermediate and final accuracy as in the first case (HRLs) means that is relatively low: a larger part of total failure comes from task-solving errors, and the model is more capable of generating successfully in the case of task-solving success. However, high intermediate accuracy and low final accuracy as in the second case (LRLs) indicates that translation loss accounts for a much bigger percentage of total error, going up to for spa-cat with aya-23, and for spa-mar with llama-3.1.
Note that with the exception of language pairs with English as a target, or with Telugu as a source, translation failure dominates total failure () for and of language pairs for aya-23 and llama-3.1. English is a special case: given that it dominates the internal representations of these models, generation into English may have a minimal or absent translation stage, meaning that task-solving success largely determines final success. We discuss the case of Telugu as a source below.
The above observations indicate that translation loss plays an outsized role in final failure even for supported target languages, and an even bigger role for low-resource or unsupported languages. These results validate the translation barrier hypothesis for the word translation task.
Evidence for entanglement in the task-solving stage
The above discussion generally supports our understanding of target-language-agnostic task-solving stage followed by a target-language-specific translation stage. However, we note the caveat that there is a clear disparity between intermediate task accuracy across target languages, with the largest difference of percentage points between spa-eng and spa-amh with aya-23, as well as differences between supported languages ( percentage point difference between hin-eng and hin-jpn with aya-23). This indicates some level of entanglement between task-solving success and language of generation, meaning that the intended language of generation affects the model’s ability to task-solve to some extent. This hints that given the current state of LLMs, even in the hypothetical case of no translation loss, end-to-end processing for low-resource or unsupported languages would still fall behind that for HRLs due to imperfect transfer of task-solving abilities to LRLs (including some supported languages). This finding calls for an evaluation of the future of end-to-end multilingual generation with LLMs on a larger scale of languages, with one alternative being to invest in better explicit output-side LLM+MT cascades.
Impact of source language
We see that for both models, mean intermediate and final accuracy as well as mean are highest for Spanish, then Hindi, then Telugu (Table 1). It is unsurprising that intermediate accuracy and therefore final accuracy depend on the source language: task-solving success is naturally conditioned on comprehension of the task inputs.999While we do not experiment with different prompt languages other than English, we can imagine that the prompt language may also show a similar effect. These findings justify exploring an input-side MT+LLM cascade for different language task inputs even for supported languages.
This also explains why is lowest for Telugu: intuitively, with low intermediate accuracy, translation loss can only occur on very few inputs as a percentage of the total failure, which is dominated by failed task-solving cases. For example, since task-solving fails for of total inputs for Telugu-Cebuano with aya-23, task-solving failure dominates total failure even though the translation stage fails to convert nearly all correct intermediate outputs to correct final outputs.
4.4 Effect of scale
We repeat our experiments for language pairs on Llama-3.1-70B-Instruct and show the effect of using a larger model on in Figure 6. We find that the model shows similar intermediate accuracy on average as the model, but consistently higher final accuracy, or capability to produce the target language, resulting in lower . However, absolute values remain high. These preliminary results indicate that the translation barrier may continue to be a problem at scale. We invite future work in better understanding the translation barrier at different scales and for other tasks.
5 Related Work
Behavior of Multilingual LLMs
While multilingual LLMs are trained to perform well for several languages (Lin et al., 2022; Workshop et al., 2023; Aryabumi et al., 2024), their pre-training data is dominated by English (Xue et al., 2021; Chung et al., 2023; Li et al., 2025a). Etxaniz et al. (2024); Zhao et al. (2024) show that the explicit cascade, i.e. translating inputs into English before passing them into LLMs, often outperforms end-to-end inference. Further, works such as Wendler et al. (2024) and (Schut et al., 2025) apply logit lens to intermediate layers and observe that multilingual LLMs “think” in English, or that intermediate representation spaces of these models lie close to English. This is reinforced by works studying neuron activations for multilingual processing (Foroutan et al., 2022; Tang et al., 2024; Kojima et al., 2024), which find language-specific neurons only in later model layers. Our work draws from this intuition with the goal of characterizing the implicit pipeline for multilingual generation and diagnosing its failure.
Cross-lingual Representations
Parallel sentences exhibit highly similar hidden states in early layers of multilingual LLMs (Muller et al., 2021; Conneau et al., 2020), and fine-tuning solely on English tasks further strengthens this alignment (Li and Murray, 2023), enabling cross-lingual transfer. Moreover, the degree of representation similarity predicts performance in other languages (Zhang et al., 2023; Li et al., 2025b; Liu and Niehues, 2025; Li et al., 2024). Our work adds to this discourse by characterizing the task-solving languages of the model, and providing evidence of some degree of target-language agnosticity of the task-solving stage.
6 Conclusion
Our paper seeks to investigate the reasons for the failure of multilingual generation with LLMs. We track the language and accuracy of intermediate model layer outputs to demonstrate a task-solvingtranslation pipeline for multilingual generation. We introduce the translation barrier hypothesis, calling attention to translation failure as an important culprit for final failure, and validate it for a word translation task by comparing task-solving as well as final success over language pairs. Our findings raise important questions for future strategies aimed at improving multilinguality in LLMs.
Limitations
Scope
The main limitation of this work is that we only explore and validate the translation barrier hypothesis for a word translation task. It may be more difficult to assess task-solving success for more complex tasks such as sentence-level machine translation, summarization, or open-ended instruction following in a principled manner using artifacts such as intermediate layer representations, since the nature of correct intermediate answer concepts could be less clear to ascertain. There may also be additional complexities to take into account when extending our setup to these tasks: for example, whether task-solving occurs “globally” over many parts or subtasks of a single input or “locally” for different fragments of the input, or the manner in which the target language syntax interacts with the ordering of the concepts in the task-solving stage. Thus, our hypothesis may be difficult to verify for other tasks.
Further, we clarify that our findings regarding the role played by the translation barrier may be dependent on the nature and complexity of the task, as well as the usage of strategies such as task-specific finetuning or inference-time techniques such as in-context learning to boost multilingual performance.
We do not claim that the statement of the hypothesis holds for all tasks and settings; however, we believe that the line of investigation in our work has important implications for future directions to improve the multilingual capabilities in LLMs, specifically regarding the potential of end-to-end versus cascaded approaches for the same. Our goal is to invite discussion and further investigation, including potential counter-evidence from other tasks, regarding the most viable strategies for high quality generation in mid- to low-resource languages with LLMs.
Models and languages
We conduct our experiments on two mainstream multilingual models; our findings should be verified on a range of different models. Similar to previous work whose insights we build upon, we also only work with decoder-only models. Our posited internal pipeline may look different for encoder-decoder models such as Aya-101 (Üstün et al., 2024).
Further, our findings are restricted to the target languages and source languages that we test on, and may differ for other languages and language families. In this work, we focus on mid- to low-resource range languages, including national languages like Swahili and Thai. Future work may look at extending our setup and exploring the translation barrier for lower-resourced languages.
Further, we use English as our prompt language regardless of the source language (i.e. the language of the term to be translated); this is as per work that says that this performs better (Dey et al., 2024). Testing the effect of different prompt languages may add further nuance to our findings.
Ethics Statement
We do not anticipate any negative consequences of this work, which is intended to diagnose poor multilingual capabilities in large language models. However, we acknowledge that different language communities may have various and diverse needs from language technologies that may diverge from the points of focus of mainstream NLP (Bird, 2020). All the models we used are publicly available for research purposes: Aya-23-8B has a CC-BY-NC license, and the Llama family of models have a Llama 3.1 Community License Agreement, permitting research use.
Acknowledgments
We would like to thank Kaiser Sun and Krithika Ramesh for their helpful feedback. This work is partially supported by the AI and Advanced Computing Institute at Schmidt Sciences.
Appendix A Languages
See Table 2 for the list of languages that we studied.
| Language | Language (Sub)Family | FLORES+ code | Supported by Aya | Supported by Llama |
| English | Germanic | eng_Latn | ✓ | ✓ |
| Spanish | Romance | spa_Latn | ✓ | ✓ |
| French | Romance | fra_Latn | ✓ | ✓ |
| German | Germanic | deu_Latn | ✓ | ✓ |
| Russian | East Slavic | rus_Cyrl | ✓ | - |
| Japanese | Japonic | jpn_Jpan | ✓ | - |
| Portuguese | Romance | por_Latn | ✓ | ✓ |
| Italian | Romance | ita_Latn | ✓ | ✓ |
| Turkish | Turkic | tur_Latn | ✓ | - |
| Korean | Koreanic | kor_Hang | ✓ | - |
| Arabic | Semitic | arb_Arab | ✓ | - |
| Polish | West Slavic | pol_Latn | ✓ | - |
| Hindi | Indic | hin_Deva | ✓ | ✓ |
| Hebrew | Semitic | heb_Hebr | ✓ | - |
| Ukranian | East Slavic | ukr_Cyrl | ✓ | - |
| Chinese (Simplified) | Sinitic | zho_Hans | ✓ | - |
| Chinese (Traditional) | Sinitic | zho_Hant | ✓ | - |
| Vietnamese | Austroasiatic | vie_Latn | ✓ | - |
| Indonesian | Austronesian | ind_Latn | ✓ | - |
| Romanian | Romance | ron_Latn | ✓ | - |
| Dutch | Germanic | nld_Latn | ✓ | - |
| Czech | West Slavic | ces_Latn | ✓ | - |
| Greek | Hellenic | ell_Grek | ✓ | - |
| Marathi | Indic | mar_Deva | - | - |
| Swahili | Niger-Congo | swh_Latn | - | - |
| Nepali | Indo-Aryan | nep_Deva | - | - |
| Tamil | Southern Dravidian | tam_Taml | - | - |
| Telugu | Southern Dravidian | tel_Telu | - | - |
| Uzbek | Turkic | uzn_Latn | - | - |
| Catalan | Romance | cat_Latn | - | - |
| Bosnian | South Slavic | bos_Latn | - | - |
| Cebuano | Austronesian | ceb_Latn | - | - |
| Persian | Indo-Iranian | pes_Arab | ✓ | - |
| Thai | Kra-Dai | tha_Thai | - | ✓ |
| Amharic | Semitic | amh_Ethi | - | - |
| Yoruba | Niger-Congo | yor_Latn | - | - |
Appendix B Dataset Details
| Language | Average Number of Equivalents in Lexica |
| English | 7.5 |
| Spanish | 9.5 |
| French | 7.7 |
| German | 9.8 |
| Russian | 12.9 |
| Japanese | 6.0 |
| Portuguese | 8.6 |
| Italian | 7.4 |
| Turkish | 9.7 |
| Korean | 8.6 |
| Arabic | 6.8 |
| Polish | 10.3 |
| Hindi | 7.0 |
| Hebrew | 3.9 |
| Ukranian | 7.4 |
| Chinese (Simplified) | 1 |
| Chinese (Traditional) | 1 |
| Vietnamese | 6.8 |
| Indonesian | 9.0 |
| Romanian | 12.9 |
| Dutch | 8.2 |
| Czech | 6.5 |
| Greek | 2.7 |
| Marathi | 2.3 |
| Swahili | 1.7 |
| Nepali | 1.1 |
| Tamil | 1.9 |
| Telugu | 1 |
| Uzbek | 4.4 |
| Catalan | 1.4 |
| Bosnian | 5.5 |
| Cebuano | 1.3 |
| Persian | 5.4 |
| Thai | 9.4 |
| Amharic | 1 |
| Yoruba | 1.4 |
In order to curate the source concepts that we translate to every language, we prompted ChatGPT to give us instances each of English nouns, verbs, adjectives and adverbs. We then filtered out any words that did not appear in the English NLTK WordNet Bird and Loper (2004); Miller (1994), and manually vetted the remaining words to retain “general” concepts (i.e. excluding any proper nouns as well as highly specific cultural terms). We also checked for offensive words and did not encounter any. Finally, we sample words per part-of-speech of the filtered words, and conduct all our experiments on the resulting concepts.
We collect translations, including synonyms and alternate word senses, for these concepts for each of the other languages studied using Google Translate as mentioned in § 2. For average number of equivalent words per source word per language, see Table 3.
Manual evaluation
To evaluate the expanded alternative translation outputs, we sampled 100 words from two languages (Turkish and Polish) and manually evaluated the alternative translation groups for each source word for plausibility (on a scale of 1 to 4, 1 = “none of the words are plausible” and 4 = “all of the words are plausible”) and for coverage (on a scale of 1 to 4, 1 = “none of the word senses are covered” and 4 = “all of the word senses are covered”). We report 3.9 and 3.4 for coverage, and 3.8 and 3.7 for plausibility, for Polish and Turkish respectively.
The Turkish annotation was performed by an author on the paper, and the Polish annotation by a volunteer (acknowledged in the final version). Both annotators had native or near-native fluency. The consent of the annotators was obtained regarding the usage of the collected data for accuracy and coverage reporting of the lexica. We did not provide any monetary compensation for this annotation.
Appendix C Model hyperparameters and prompting setup
Hyperparameters and compute
We use greedy decoding with a temperature of and the official chat template of aya-23 and llama-3.1, respectively. We run inference with logit lens on NVIDIA H100 NVL Tensor Core GPUs. This takes about minutes to process the test set of size with batch size on a single GPU, for each of our language pairs for each of aya-23 and llama-3.1. We downloaded aya-23 and llama-3.1 from Huggingface Transformers (Wolf et al., 2020).
We additionally run experiments on language pairs with llama-3.1-70B. This takes minutes per language pair on GPUs with a batch size of .
This does not include exploratory experiments and debugging.
Prompting
Given that LLM performance is sensitive to the wording of the prompt (Anagnostidis and Bulian, 2024), we try a few different prompts for the word translation task. We find that final performance differs slightly with different wordings of the instruction as well as the manner in which the target language is specified.
Here are examples of prompts we explored:
1.
Translate the following word into English:
gata
2.
Translate the following word from spa_Latn to mar_Deva. Respond with a single word.
Word: gata
Translation:
3.
Give a one-word translation of the following word from Spanish to Marathi.
Word: gata
Translation:
4.
Translate the following word from Spanish to Marathi. Respond with a single word.
Word: gata
Translation:
We use prompt (4) for all our experiments, as performing well for both our models.
Appendix D Examples of inputs and outputs
See Table 4 for examples of intermediate layer outputs for some inputs.
Appendix E Reliability of the LangID
Recent work shows that neural language identification models can be performant even for mixed-script and noisy multilingual data Sirin et al. (2024). To evaluate the language ID performance of the NLLB model NLLB Team et al. (2022) over intermediate outputs, we sampled source-target word pairs from two different language pairs, Hindi - Turkish and Spanish - German. Note that we discard labels that do not correspond to one of our target languages as per § 3, and only report on-targetness or off-targetness in Figure 2 for inputs for which we can find a reliable label.
We evaluated the resulting language identification labels for two different layer outputs for each word pair. We found that all labeled samples had outputs that were recognizably in some language. We marked the LID as correct if it accurately identified the language of any of the words in the output, given that the outputs could be language-mixed. We found that the model was accurate over intermediate layer outputs.
Appendix F Evolution through layers for all languages
Demonstrating abstract task-solving
We show the percentage of final answers that were identified as correct in some intermediate layer, in a different language from the target, in Figure 7 for both models. In general, we see that this percentage is high, although less so for HRL supported languages. This, in conjunction with the shape of the distribution over intermediate languages that are largely target-language agnostic as discussed in § 4 and Appendix H, demonstrates that intermediate layers perform task-solving at some layer of abstraction.
Layer-wise accuracy for other language pairs
We choose five representative language pairs and show the evolution of on-target and off-target accuracy through various layers for each of them for aya-23 and llama-3.1 in Figure 8. These cover all three source languages, supported HRL and MRL languages, as well as unsupported languages.
Appendix G Layer of switch for individual language pairs
In Figure 10, we show the layer with the maximum increase in target language presence for accurate outputs, in order to identify a particular layer of switch between the task-solving and translation stages. We exclude language pairs that have negligible final accuracy (), since these effectively have no translation stage. We see the layer of switch is on or after the fourth-last layer for aya-23 and on or after the third-last layer for llama-3.1, with aya-23 often showing earlier switches than llama-3.1.
Appendix H Distribution of “task-solving” languages by target language
Given that there is a prior model-internal reasoning step before generation, we are interested in the language(s) in which this reasoning occurs. As we visualize in Figure 2 and Figure 8, there is a somewhat fluid shift to the target language at later layers; we can consider these the layers responsible for the translation / generation part of the pipeline. We see that this shift emerges around the fourth-last layer; we therefore consider the task-solving layers as all the layers up to the fourth-last layer. We show the distribution of correct intermediate equivalents over intermediate languages for each target language (averaged over source language) in Figure 11 for llama-3.1 for these middle layers. In general, we see that this distribution looks similar over target languages, supporting the idea of model interlingua that is to a large degree agnostic of the target language. The plot looks similar for aya-23 and also largely similar when using different layer cutoffs for what we consider as task-solving layers, with increasing target language presence as we move closer to the final layer.
Appendix I On-targetness
We show the percentage of final layer on-target outputs in Figure 12 for both models. Predictably, this is higher for supported HRLs than for unsupported or low-resource languages. Note that these results are dependent on the accuracy of the language ID model, which may be easily confused on short spans of text.
Appendix J Results for all languages
Appendix K Examples on Translating Sentences
While this work focuses on single word translation, we show examples in this sentence of applying iterative logit lens as described in § 3 for a sentence-level translation task. We observe that intermediate layers exhibit code-switching behavior that interpolates between the correct target language and English. We invite future work in studying the translation barrier for sequence-level tasks.
K.1 Example 1
| Source Sentence | Mes 3 minutes n’ont pas déjà commencé, n’est-ce-pas ? |
| Correct Translation | {CJK*}UTF8gbsn我的三分钟还没开始吧? |
| Layer 32 (Final Layer) | {CJK*}UTF8gbsn我的3分钟还没有开始,是吗? |
| Layer 29 | {CJK*}UTF8gbsn我的3 minute还没有开始,是吗? |
| Layer 27 | my three minutes chua chua{CJK*}UTF8gbsn开始吧 right{CJK*}UTF8gbsn吗? |
| Layer 25 | my three minutes already chua{CJK*}UTF8gbsn开始吧 right{CJK*}UTF8gbsn吧? |
K.2 Example 2
| Source Sentence | Je veux dire, je suis déjà assez nerveuse comme ça. |
| Correct Translation | {CJK*}UTF8gbsn我是说,我已经够紧张的了 |
| Layer 32 (Final Layer) | {CJK*}UTF8gbsn我的意思是,我已经很紧张了。 |
| Layer 29 | {CJK*}UTF8gbsn我的意思是我我 already足够紧张了。 |
| Layer 27 | {CJK*}UTF8gbsn我mean是我我 already cukup nervous了。 |
| Layer 25 | {CJK*}UTF8gbsn我 already là, already already already nervous了。 |
K.3 Example 3
| Source Sentence | De plus, au Canada, nous avons un formidable système de santé. |
| Correct Translation | {CJK*}UTF8gbsn在加拿大,我们有很好的医保制度 |
| Layer 32 (Final Layer) | {CJK*}UTF8gbsn此外,在加拿大,我们有一个很棒的医疗系统。 |
| Layer 29 | {CJK*}UTF8gbsn此外,我们加拿大,我们拥有 fantast好的的健康 system。 |
| Layer 27 | {CJK*}UTF8gbsn此外,我们 Canada我们我们有 fantast好的的 healthcare system。 |
| Layer 25 | {CJK*}UTF8gbsn此外,我们 Canada we我们有 superb好的的 healthcare system。 |
K.4 Example 3
| Source Sentence | Mais, tout d’abord, vous devez savoir 2 choses à mon propos |
| Correct Translation | Aber Sie müssen zuerst zwei Dinge über mich wissen. |
| Layer 32 (Final Layer) | Aber zuerst, Sie müssen 2 Dinge über mich wissen. |
| Layer 29 | But zuerst müssen ihr müssen zwei2 things über mich wetenissen. |
| Layer 27 | but first einmal you must know two things things about me know know. |